### v3

# The Causal Effect of the Two-For-One Strategy in the National Basketball Association

Abstract
This study evaluates the effectiveness of the “two-for-one” strategy in basketball by applying a causal inference framework to play-by-play data from the 2018-19 and 2021-22 National Basketball Association regular seasons. Incorporating factors such as player lineup, betting odds, and player ratings, we compute the average treatment effect and find that the two-for-one strategy has a positive impact on game outcomes, suggesting it can benefit teams when employed effectively. Additionally, we investigate potential heterogeneity in the strategy’s effectiveness using the causal forest framework, with tests indicating no significant variation across different contexts. These findings offer valuable insights into the tactical advantages of the two-for-one strategy in professional basketball.
Executive Summary
The two-for-one basketball strategy involves attempting to take a shot early in a possession in order to maximize the number of possessions at the end of a period. Based on our analysis of two seasons of data from the National Basketball Association, we estimate that teams that attempt a two-for-one strategy gain slightly more than half a point per attempt, on average, compared to teams that do not attempt a two-for-one strategy. We do not find that differences in certain relevant covariates lead to substantial differences in this point gain. Our analysis thus empirically validates the growing consensus of the effectiveness of the two-for-one strategy in the National Basketball Association.
Keywords— causal inference, average treament effect, heterogenous treatment effect, sports analytics
1 Introduction
In recent years, the use of data and statistical methods to analyze sports strategies has gained significant traction, driven by the availability of detailed game data and advances in data science. In professional sports leagues like the National Basketball Association (NBA), decision-makers seek to optimize team performance through strategic adjustments informed by empirical analysis. Among these strategies, the “two-for-one” tactic, where a team attempts to gain an extra possession by timing shots near the end of a period, has become popular. However, despite its widespread use, little rigorous evidence exists on whether this strategy provides a measurable advantage. This study aims to address this gap by applying causal inference methods to determine the effectiveness of the two-for-one strategy.
An increasing body of research has applied causal inference methods to investigate the causal effects of various sports strategies. For example, Gibbs et al. (2022) utilize causal inference to evaluate the impact of calling a timeout to disrupt an opposing team’s scoring run in NBA games. Nakahara et al. (2023) analyze pitching strategies in baseball through a causal inference framework, while Yam and Lopez (2019) examine the decision-making process surrounding fourth-down plays in the NFL, aiming to quantify the effectiveness of these high-stakes decisions. Additionally, Wu et al. (2021) explore the effect of crossing the ball in soccer, seeking to determine whether this strategy significantly influences scoring opportunities.
These studies represent a growing interest in the causal analysis of strategic decisions in sports, where the interplay of chance and skill presents unique challenges to identifying causality. By employing techniques that control for confounding variables and capitalize on observational data, such research provides insights that go beyond traditional statistical descriptions, aiming instead to uncover the causal impact of specific decisions. The increasing application of causal inference in sports analytics thus reflects a broader trend towards data-driven decision-making and evidence-based optimization of strategy across various sports contexts.
Causal inference (Imbens and Rubin, 2015) methods are essential tools for rigorously analyzing the effects of treatments or interventions in observational settings, particularly when randomized controlled trials (RCTs) are impractical, as is often the case in sports. In this study, we focus on estimating the average treatment effect (ATE) of the NBA’s two-for-one strategy. The ATE quantifies the average difference in outcomes (e.g., difference in score margin) between situations where the two-for-one strategy is employed and those where it is not. Estimating the ATE in observational data, however, requires addressing confounding variables—factors that influence both the treatment and the outcome, potentially biasing the results. To mitigate such bias, we apply augmented inverse probability weighting, a widely used method in causal inference that allows for more reliable estimation of treatment effects. This method involves estimating the probability (propensity score) that a given observation will receive the treatment (the two-for-one strategy) based on observed covariates (e.g., player lineup, time in the game). Each observation is then weighted by the inverse of its propensity score to create a pseudo-population with a similar distribution of covariates between treated and control groups.
Beyond estimating the ATE, it is crucial to assess whether the treatment effect varies across different contexts or player configurations. To do this, we employ the causal forest (Athey and Wager, 2019; Wager and Athey, 2018) framework, a machine learning-based method that enables the estimation of heterogeneous treatment effects (HTE). Traditional ATE estimation assumes that the treatment effect is constant across all observations, but in real-world settings like the NBA, the effectiveness of a strategy is likely to vary based on game context, player lineups, and other factors. The causal forest framework addresses this by partitioning the data into subgroups using decision trees (Hastie et al., 2009). Each subgroup represents a distinct set of conditions where the treatment effect may differ. Causal forests construct many such trees and aggregate the resulting estimates to produce a flexible, non-parametric estimate of treatment effect heterogeneity. By identifying subgroups with distinct treatment effects, causal forests allow us to uncover whether the two-for-one strategy is more or less effective in different situations, such as when certain players are on the court or at different points in a game.
In addition to estimating treatment effect heterogeneity, we apply the concept of rank average treatment effect (RATE) (Yadlowsky et al., 2025) to better understand the strategy’s impact. RATE works by ranking observations based on their predicted treatment effects, identifying which game contexts or player configurations are most likely to benefit from the two-for-one strategy. The RATE method quantifies the expected impact on the top ranked subgroups and compares these with the average effect across all observations. This ranking approach is particularly valuable in practice, as it highlights the specific conditions under which the two-for-one strategy is most advantageous, offering actionable insights for coaches and analysts looking to optimize in-game decisions.
The structure of this paper is as follows. Section 2 presents definitions and background on the two-for-one strategy and details the data set used in our analysis. Section 3 outlines the causal methodology employed to assess the impact of this strategy. Section 4 presents the results obtained from analyzing two seasons of NBA data, highlighting key findings on the effectiveness of the two-for-one approach. Section 5 concludes with a brief discussion.
2 Definitions and Data
In this section, we provide a detailed explanation of the two-for-one (TFO) strategy in the NBA, including a mathematical definition of the strategy and a description of the data sources used in this study.
The TFO (Feng, 2015; Fischer, 2015) strategy in the NBA is a time-management tactic that aims to maximize a team’s scoring opportunities by securing two offensive possessions at the end of a period, while limiting the opponent to just one. This strategy is commonly employed in the last 30–40 seconds of a period, especially during close games, where maximizing scoring chances can be crucial. The TFO strategy is primarily focused on managing the game clock to optimize possessions and increase scoring potential before the period ends. To implement the TFO, a team aims to take a shot with a certain amount of time remaining in the period so that even if the opponent uses a full possession, the team can secure the ball again before time expires.
In order to attempt a TFO, a team may be required to take a shot early in the possession, regardless of the quality of shot. This runs counter to the common strategy in basketball of progressing through an offensive set until an opportunity for a quality shot is realized. Thus the potential down side of attempting a TFO is that even if a team ends up with more possessions at the end of the period than their opponent, they may not ultimately benefit because of poorer scoring opportunities.
2.1 Definitions
To see if the potential gains of a TFO strategy outweigh the potential losses we need to determine a causal effect, and for that we need a well-defined treatment. That is, we need to be able to classify scenarios during an NBA game as those in which a team enacted the TFO strategy, and those in which they did not. This is not a trivial matter. Teams do not announce that they are attempting a TFO, and it is unclear how attune players are to the game and shot clock during any particular play. Thus we rely on a combination of time and play outcomes in order to identify TFO opportunities. These are scenarios when a team comes into possession of the ball with enough time left in the period such that taking a shot or initiating offense relatively early in the shot clock leads to a increased chance to get an extra possession, while waiting until the end of the shot clock to do these things increases the chance the opposing team will have the last possession. We then categorize those TFO opportunities into attempts (treatment group) and non-attempts (control group).
Several important characteristics of NBA game play inform these definitions. NBA games are played in four 12-minute periods. Possession at the beginning of the game is determined by a jump ball, the result of which also determines which team gets possession at the start of subsequent periods. The shot clock in the NBA is 24 seconds, meaning that from the time a team gains possession of the ball after a turnover, defensive rebound, or made basket by the opposing team, the team has 24 seconds in which to attempt a shot. If a team rebounds their own missed shot the shot clock is reset to 14 seconds instead of 24. If a team gains possession with less than 24 seconds left in the period, the shot clock is turned off and the team in possession can attempt to have the last possession of the period by waiting until the very end to take a shot. Therefore, if a team wants to improve its chances of gaining the last possession of the period, it should hope that the opposing team gains possession with more than 24 seconds left.
However, it can take up to a few seconds for possession to change, particularly after a missed shot. As an example, suppose a player on Team A takes a shot with 6 seconds left in the period. It takes one second for the ball to get from the player’s hand to the rim, another second to bounce on the rim a couple of times before coming to the ground, and a third second before a player on Team B possesses it, now with only 3 seconds left in the period. Furthermore, if there is too little time left in the period when a team gains possession, that team has a very small chance of scoring any points.
With these characteristics in mind, we set a limit of 28 seconds left in the period, and say that if a team takes a shot after this point and the possession changes, it is unlikely that they will get a quality possession at the end of the period. Working backwards, then, a team needs to gain possession of the ball with enough time before the 28 seconds mark in order to potentially attempt a reasonable shot by this point. We set this lower limit to be 35 seconds. We choose 43 seconds as an upper limit on when a team can gain possession to exclude those scenarios where a team can get to this optimal shot time simply by waiting until the shot clock for their possession is close to 0. Thus we define a TFO opportunity as one in which a team comes into possession of the ball with between 35 and 43 seconds left in the period.
If a team presented with a TFO opportunity takes a shot or is fouled while taking a shot with at least 28 seconds to go in the period, we consider that a clear TFO attempt. Other play results are harder to categorize. We also consider an opportunity an attempt if the team is fouled with at least 28 seconds to go in the period, with the rationale being that fouls more often occur when a team is trying to initiate offense. On the other hand, we consider an opportunity to be a non-attempt if the team maintains possession until there are less than 28 seconds left in the period without taking a shot or being fouled.
We remove all other play outcomes that occur during a TFO opportunity from consideration rather than defining them as either an attempt or a non-attempt. The most common of these is a turnover. Turnovers can happen while a team is initiating offense, such as a charge foul, or just trying to hold the ball until the end of the shot clock, such as a steal that happens far away from the basket. Due to the lack of information on turnovers available in play-by-play data, we choose to remove them from our analysis. The same is true for other, less common plays such as a jump ball, a defensive three-second violation, and a kicked ball violation. Figure 1 provides a visual description of our treatment and control groups.
Figure 1: Diagram of the treatment definition.
We believe this to be a reasonable treatment definition for the TFO strategy. However, it also has limitations. One of these is that the timing cutoffs, while well motivated, are nevertheless somewhat arbitrary. Thus in Section 4.1.4 we present a sensitivity analysis in which we adjust these cutoffs slightly to explore the robustness of our estimates to changes in the definition.
A second limitation is that, as mentioned earlier, there is no way to know the intent of the coaches or the players on the floor with respect to the TFO strategy, which forces us to rely on event data for our treatment definition. There are certainly scenarios when a player intends to implement a TFO strategy but is not part of our treatment group because, for example, they lost control of the ball while going up for a shot and it is recorded as a turnover. The distinction from the clinical trials literature between intent-to-treat (ITT) and per-protocol (PP) analyses is relevant here. In the former case subjects are analyzed based on the original intended treatment, where in the latter they are analyzed based on the treatment they actually receive. Molero-Calafell et al. (2024) discuss the strengths and limitations of these analyses in relation to whether a study is more explanatory, meaning executed under more ideal experimental conditions, or pragmatic, executed under more real world conditions. Certainly our study is more pragmatic by their definitions, and thus they would recommend an ITT analysis with perhaps a PP analysis as a complement. However, in this framework we do not know the original intent and thus our approach is more similar to that of a PP analysis because an ITT analysis is not possible. However, Molero-Calafell et al. (2024) mention that employing advanced statistical methods such as the inverse probability weighting we use in our analysis can help enhance the validity of PP estimates.
We define our response variable as the Post Opportunity Difference (POD), which measures the difference in score differential from the time of the TFO opportunity to the end of the period. A positive POD indicates that the team increased its score margin by the end of the period. Conversely, a negative POD implies that the team’s score margin decreased after the TFO opportunity. Thus if the TFO strategy is effective the POD would be higher for teams that attempt a TFO than those who do not.
2.2 Data
The main source of data we use is play-by-play data from the NBA. This is a list of all the different plays that took place in an NBA game as well as supplemental information such as the players involved in the play, the score of the game at the time of the play, the period the play took place in, and the amount of time left in that period. We obtain play-by-play data from the nbastatR package (Bresler, 2024), which interfaces with the NBA Stats API and other reference sources to give a variety of data and summary statistics at the player, team, game, and season level. This includes play-by-play data dating back to the 1996-1997 season, the first year the NBA kept such data. For our analysis we focus on two recent seasons, that of 2018-2019 and 2021-2022. The 2018-2019 season was the first regular season under a modified set of shot clock rules as well as the last regular season before the Covid-19 pandemic altered the NBA schedule, and the 2021-2022 season was the first full season after the pandemic, making these natural seasons to consider in order to explore the overall effect of the TFO strategy as well as possible heterogeneity across multiple seasons.
Using this play-by-play data, we identify all TFO opportunitys in the first three periods of each NBA regular season game. We choose to exclude the fourth periods and overtime because the behavior of NBA games is often much different at the end of those periods. If the game is not close, teams are no longer necessarily worried about maximizing their possessions because of the feeling that nothing they do will affect the outcome. If the game is very close, teams tend to utilize a strategy of intentionally fouling in order to increase the number of possessions in the game. We believe that excluding these scenarios allows us to better isolate the treatment, and we present a numerical justification for this choice in Section 4.1.5. We also exclude TFO opportunities in which the team has already had an earlier TFO opportunity in the same period. For the 2018-2019 season, there were a total of 1567 TFO opportunities, of which 1036 were classified as attempts and the remaining 529 classified as non-attempts. For the 2021-2022 season there were a total of 1479 TFO opportunities, of which 950 were classified as attempts and the remaining 462 classified as non-attempts. Each NBA season contains 1230 regular season games, so there was an average between 1 and 2 TFO opportunities per game based on our criteria.
To enhance our analysis with relevant covariates, we draw from several supplemental data sources. First, to gauge the relative strength of the teams at the time of each game, we incorporate betting odds, which indicate the favored team and the expected point spread. These odds effectively summarize expert assessments of each team’s strength. Additionally, to measure the projected pace of play, we include betting totals, which provide a prediction of the total points expected to be scored in a game. Both the betting odds and totals are sourced from the Sportsbook Reviews archive (Sportsbook Reviews, 2023).
For player-specific information, we use individual player ratings from the NBA 2K video game series (HoopsHype, 2018, 2021) which provide relative strength assessments for each player. While not perfect, these ratings are a reasonable proxy for the overall ability of players. In addition to previous years’ statistics, ratings are informed by expert projections, particularly for players new to the league (Irving, 2021; Panerio, 2024). We obtain starting lineups from BasketballReference.com (Basketball Reference, ) and use this information, combined with play-by-play data, to determine which players are on the court at any given time. These covariates allow us to account for variations in team and player strength as well as game tempo, enhancing the robustness of our analysis. Table 1 provides the name and definitions of the variables we use.
Variable Name Variable Definition
Period Period of the game; data is restricted to the first three periods.
Time Left Remaining time in the current period, measured in seconds.
Score Margin Difference between the scores of the two teams.
Max Rating Maximum player rating among players on the court for the team at that moment.
Max Rating Opposition Maximum player rating among players on the court for the opposing team at that moment.
Mean Rating Average player rating of the players on the court for the team at that moment.
Mean Rating Opposition Average player rating of the players on the court for the opposing team at that moment.
Spread The projected difference in final score for the two teams, set by oddsmakers.
Total Score The projected sum of the scores of the two teams, set by oddsmakers.
Table 1: List of variables used in the propensity score model along with their definitions.
Combining these covariates and the play-by-play data into a single usable data set requires a substantial amount of data cleaning. The play-by-play data includes substitutions but not a list of the current players on the floor, so we have to input the starting lineups and then create variables that track the current lineup on the floor. Player names are also represented differently in the various sources which complicates data merging. Finally, the play-by-play data provides narrative descriptions such as the type of shot, which is a challenge when categorizing plays. As an example, in Table 2 described in the section below, the phrases jump shot, finger roll layup, and dunk all indicate a shot has been attempted.
2.3 Example
To illustrate the definitions and data described above, consider an example from a game played between the Phoenix Suns and the Golden State Warriors on October 22, 2018. Table 2 shows several rows of the play-by-play data with a subset of relevant columns. In the first period of the game the Suns gained possession with 42 seconds left after a Warriors turnover, giving the Suns a TFO opportunity. However, they did not take a shot early in the shot clock. Rather, they committed a turnover with 25 seconds left. By our definition, this is a non-attempt. The Warriors gained possession with the turnover and made a 3 point shot, after which there was no more scoring in the period. So from the time of the TFO opportunity until the end of the period the Suns scored 0 points and the Warriors scored 3 points, for a POD of -3.
With 37 seconds left in the second period of the same game, the Warriors gained possession after a made shot by the Suns, giving the Warriors a TFO opportunity. With 31 seconds left Kevin Durant made a dunk for the Warriors, and we consider this shot attempt with more than 28 seconds to go in the period a TFO attempt. The Suns gained possession at that point and committed a turnover with 12 seconds left. Thus the Warriors gained a final possession, and Quinn Cook made a three-point shot with one second left, which did not allow the Suns to take another shot. From the time of the TFO opportunity until the end of the period, the Warriors scored 5 points and the Suns scored 0 points, for a POD of 5.
Pr. Time DescriptionPlayHome DescriptionPlayVisitor ScoreH ScoreV
1 0:42 Curry Out of Bounds - NA NA NA
Bad Pass Turnover
Turnover (P3.T5)
1 0:25 NA Booker Discontinue NA NA
Dribble Turnover (P2.T3)
1 0:25 SUB: Cook FOR Looney NA NA NA
1 0:25 SUB: Thompson FOR NA NA NA
McKinnie
1 0:07 Jerebko 27’ 3PT Step NA 32 23
Back Jump Shot (3 PTS)
(Thompson 2 AST)
1 0:03 Iguodala STEAL (1 STL) Booker Lost Ball Turnover NA NA
1 0:01 MISS Curry 55’ 3PT NA NA NA
Jump Shot
1 0:$0$ WARRIORS Rebound NA NA NA
2 0:37 NA Booker 4’ Driving Finger 65 47
Roll Layup (14 PTS)
2 0:31 Durant 3’ Driving Dunk NA 67 47
(15 PTS)
2 0:12 NA Booker Out of Bounds NA NA
Lost Ball Turnover (P5.T9)
2 0:12 SUB: Cook FOR Jones NA NA NA
2 0:12 NA SUB: Bridges FOR NA NA
Anderson
2 0:01 Cook 24’ 3PT Pullup NA 70 47
Jump Shot (4 PTS)
(Thompson 3 AST)
Table 2: Section of play-by-play data obtained from the nbastatR package. The variables shown are the period of play, the time left in that period, descriptions of the play for both the home and visitor teams, and score for both the home and visitor teams. Variable names have been altered.
3 Methodology
Using the definitions from the previous section, we define the treatment $W$ for TFO opportunity $i$ as $W_{i}=1$ if the TFO is attempted and $W_{i}=0$ if the TFO is not attempted. We let $Y_{i}$ be the POD. We employ the potential outcomes framework (Splawa-Neyman, 1923; Rubin, 1974) in order to define the primary causal effect of interest. Let $Y_{i}(1)$, $Y_{i}(0)$ be the potential outcomes under treatment and control, respectively. That is, $Y_{i}(1)$ is the POD a team would get if it attempts a TFO at opportunity $i$, and $Y_{i}(0)$ is the POD a team would get if it does not attempt a TFO at opportunity $i$. A natural estimand of interest is the average treatment effect (ATE):
$ATE=E[Y_{i}(1)-Y_{i}(0)].$ (1)
The ATE quantifies how much larger the POD would be, on average, if NBA teams presented with a TFO opportunity always made a TFO attempt rather than never made a TFO attempt.
The fundamental problem of causal inference is that we do not observe both $Y_{i}(1)$ and $Y_{i}(0)$ for any particular observation. If the scenarios where a team attempted a TFO were the same, on average, in all meaningful ways to the scenarios where a team did not attempt a TFO, an assumption known as exchangeability, then we could estimate the ATE from the data as $\hat{E}(Y_{i}|W_{i}=1)-\hat{E}(Y_{i}|W_{i}=0)$. Unfortunately that assumption is not reasonable in this situation, as potentially confounding variables such as the score margin and the time left in the period are likely to influence a team’s decision.
The methods of causal inference have been developed for precisely such circumstances, when a causal effect estimate is desired for observational data where exchangeability does not hold. Many of these methods rely on a common set of assumptions, namely consistency, no interference, conditional exchangeability, and positivity. Below we discuss these assumptions and why they are reasonably satisfied for our data, followed by a description of the specific methods we will be using to estimate treatment effects.
3.1 Causal Assumptions
The consistency assumption implies that there is a well-defined treatment with no hidden variations. In our setting we recognize that not all TFO attempts look the same, as highlighted in the previous section. Some end in a shot while others end in a foul. In some cases the team that had the TFO attempt gets the ball back, and in other cases it does not. However, because these differences are a part of the definition of our treatment, they are not hidden and do not form an egregious violation of the consistency assumption.
The no interference assumption says that the outcome for a particular unit does not depend on the treatment assignment of any other unit. Mathematically, this means that for any individual $i$,
$Y_{i}(W_{i},W_{-i})=Y_{i}(W_{i})$ (2)
where $W_{-i}$ represents the treatment vector excluding $i$. This assumption would be violated in our setting if the success of a particular team’s TFO attempt was affected by whether another TFO opportunity in the study was an attempt or a non-attempt. One potential source of interference could come from the possibility of multiple TFO opportunities in the same period. For example, suppose a team gains possession with 43 seconds left and makes a shot with 36 seconds left, after which the opposing team gains possession and makes a shot with 29 seconds left. Then both teams have attempted a TFO. However, we do not believe that the success of the TFO attempt by the team gaining possession with 36 seconds left in this scenario depends on the fact that the first team is also attempting a TFO. Had the first team gained possession with 50 seconds left instead of 43, we would not consider this a TFO attempt but the second team gets the ball in an practically identical situation as the original scenario. Thus we have left these instances in our data set.
A slightly more complicated scenario occurs if the same team has back-to-back TFO opportunities. This may happen, for example, if a team gains possession with 43 seconds left, misses a shot with 36 seconds left, but gets an offensive rebound, thus earning another possession. Often in this case the player that gets the rebound is able to take a quick shot close to the goal which has a higher than average probability of going in or drawing a foul. So it is plausible that these TFO attempts may lead to a higher POD because they follow another TFO attempt from the same team. Thus we have removed these instances from our data set.
Other than units occuring in the same period, it is difficult to imagine a scenario where there would be a serious violation of this assumption. It is possible due to the season-long nature of the data that if a particular team has a large number of TFO attempts relative to non-attempts early in the year, the players may practice last-second shots more because they have been experienced them more in the course of play, and thus they could improve the outcome of the TFO attempts later in the year. However, players already practice these types of shots because of their importance to game play, and it seems unlikely that the amount of time they spend practicing such shots substantially depends on how often they attempt a TFO strategy. Thus the no interference assumption seems to reasonably hold for our setting.
Conditional exchangeability implies that, given a set of covariates $X$, the treatment assignment $W$ is independent of the potential outcomes:
$Y_{i}(0),Y_{i}(1)\perp W_{i}\mid X_{i}.$ (3)
This allows us to assume that treated and untreated individuals with the same values of $X$ have comparable distributions of potential outcomes, thus enabling causal comparisons. Suppose our set of covariates consists of the score margin and the time left in the period. Then conditional exchangeability would say, for example, that TFO opportunities where a team is down by 4 with 40 seconds left in the period are the same, on average, for both the set of attempts and non-attempts.
One way to examine the reasonableness of this assumption is to model the relationships between treatment, response, and covariates in a directed acyclic graph (DAG). In these diagrams, an arrow from one node to another represents a causal relationship between those variables. For a DAG to be complete, it must contain all common causes of any variables in the graph. Under this assumption, we can use a quick visual check called the back-door criteria to determine whether a set of covariates satisfies the conditional exchangeability assumption. See Pearl (2009) for more information on the use of DAGs in causal inference.
We propose the DAG given in Figure 2 as a model for our causal structure, using the covariates in Table 1. Note that we combine the ratings variables into a single node for the sake of simplicity and because we think they would have the same relationship structure with the other variables. We think that each of these covariates could reasonably affect both the treatment and response variables. Teams may be more likely to attempt a TFO if there is more time left in the period, and this also could possibly allow for more possessions after the TFO opportunity, which could increase the POD. There spread and the score margin represent the expected and current point difference between the teams, and it is possible that teams that are behind or are expected to be behind may feel more of a need to maximize possessions. At the same time these teams may score less on average per possession than those winning or expected to win, affecting the POD. The total score is in some ways a proxy for pace of play. Teams that play faster may be more likely to attempt a TFO than a more methodical team, and a faster pace can lead to more possessions and thus affect the POD. The quality of the players on the floor certainly impacts the scoring at the end of the period, but it may also impact the likelihood of a TFO attempt if, for example, a team playing its best players is able to execute well against an opponent playing weaker players. Finally, we allow for the existence of period specific effects as team strategies may well be different during different parts of the game.
Figure 2: DAG of our proposed causal structure, created using the DAGitty web application (Textor et al., 2017)
In addition to these confounding relationships, there are several relationships between covariates included in the DAG. For example, some coaches have set substitution patterns, and so the period of the game affects the ratings of the players on the floor. Alternatively, if a team is ahead by a large amount, a coach may change this pattern and rest his best players more than in a close game. It is important that none of these extra relationships lead to a cycle in our graph, which is not allowed in a DAG and would violate temporal restrictions of causality. Letting $X$ be the set of covariates listed in our DAG, the back-door criteria and thus conditional exchangeability is satisfied. However, while we believe that we have included the most important sources of confounding, there is the possibility of unmeasured confounding in any causal analysis. In Section 4.1.3 we examine the sensitivity of our estimates to unmeasured confounding. It is also appropriate to check conditional exchangeability by looking at the balance of the covariates in the groups created by the methods discussed below, which we do in Section 4.1.1.
The final assumption necessary is the positivity assumption, which says that each TFO opportunity must have a positive probability, conditional on our set of covariates, of being in either the treatment or control group i.e.,
$0<P(W_{i}=w_{i}\mid X_{i})<1,\quad w_{i}=0,1$ (4)
There is a known trade-off between positivity and conditional exchangeability. If you condition on too many covariates, you may satisfy the latter but not the former. If you do not condition on enough covariates, the opposite may be true. We feel that we have found a good balance between the two by carefully selecting a small but rich set of covariates. We perform a visual check of the positivity assumption in Section 4.1.
3.2 Causal Methods for Average Treatment Effects
Many causal inference methods rely on the concept of a propensity score (Rosenbaum and Rubin, 1983), which is defined as the probability of receiving treatment given baseline covariates $X$. For a binary treatment $W\in\{0,1\}$, the propensity score denoted by $e(X)$ is given by
$e(X)=P(W=1\mid X).$ (5)
Propensity scores are generally unknown quantities but can be estimated using logistic regression or machine learning methods.
Propensity scores are used to estimate treatment effects in several ways. Inverse probability weighting (IPW) is a class of estimators that assigns weights to individuals to create a pseudo-population where the distribution of covariates is balanced between treatment groups. These weights are larger for individuals in groups with a lower probability of treatment assignment, adjusting for selection bias by up-weighting underrepresented individuals (Hirano et al., 2003). If conditional exchangeability holds in the true population, then marginal exchangeability holds in the pseudo-population. Thus the ATE can then be estimated by the weighted difference in outcomes,
$\hat{ATE}_{IPW}=\frac{1}{n}\sum_{i=1}^{n}\Big(\frac{W_{i}Y_{i}}{\hat{e}(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-\hat{e}(X_{i})}\Big),$ (6)
where $\hat{e}(X_{i})$ is the estimated propensity score.
Note that the set of covariates $X_{i}$ contains information about both the treatment assignment and the outcome. This motivates an alternative approach known as augmented inverse probability weighting (AIPW) (Glynn and Quinn, 2010). In addition to estimating propensity scores, this method estimates the outcome under both treatment and control, denoted $\hat{E}(Y_{i}|W_{i}=1,X_{i})$ and $\hat{E}(Y_{i}|W_{i}=0,X_{i})$, respectively. The IPW estimator is then adjusted by a term involving the weighted average of these two estimators:
$\hat{ATE}_{AIPW}=\frac{1}{n}\sum_{i=1}^{n}\left\{\Big(\frac{W_{i}Y_{i}}{\hat{e}(X_{i})}-\frac{(1-W_{i})Y_{i}}{1-\hat{e}(X_{i})}\Big)-\frac{W_{i}-\hat{e}(X_{i})}{\hat{e}(X_{i})(1-\hat{e}(X_{i}))}\right.\\ \left.[1-\hat{e}(X_{i})\hat{E}(Y_{i}|W_{i}=1,X_{i})+\hat{e}(X_{i})\hat{E}(Y_{i}|W_{i}=0,X_{i})]\right\}.$ (7)
This AIPW estimator is doubly-robust in that it is consistent for the ATE whenever either the propensity score model or the outcome model is correctly specified. It also requires the conditional exchangeability assumption mentioned above.
3.3 Causal Methods for Heterogenous Treatment Effects
While the ATE summarizes the overall impact of executing a TFO, it is natural to expect that its effect varies across different possession contexts. For example, some teams are naturally more effective at generating quick, quality scoring chances under time pressure, whereas others tend to rely on slower, more structured offensive sets that may yield lower-quality shots when rushed. Likewise, a TFO taken when a team holds a comfortable lead may influence subsequent decision-making differently than one taken in a highly competitive situation. The defensive characteristics of the opponent also matter: facing a unit that is slower to adjust to rapid possessions may create more favorable scoring conditions than facing a unit that is disciplined and well-positioned. These contextual differences may naturally lead to heterogeneous treatment effects. Using causal forests to detect this heterogeneity helps clarify when and in what types of game states the 2-for-1 is most advantageous, complementing the ATE without affecting its interpretation or underlying assumptions.
The Conditional Average Treatment Effect (CATE) (Athey and Imbens, 2016) is a statistical measure used to capture the heterogeneity in treatment effects across different subgroups within a population. It extends the concept of the ATE, which calculates the expected difference in outcomes due to a treatment across the entire population, by conditioning on specific covariates $X$. This approach is particularly useful in causal inference to identify how treatment efficacy varies across different individuals or groups, allowing for more personalized interventions.
Mathematically, the CATE is defined as:
$CATE=E[Y_{i}(1)-Y_{i}(0)\mid X=x].$ (8)
By conditioning on $X=x$, CATE captures the expected treatment effect for individuals with characteristics $X=x$. Assuming conditional exchangeability and consistency, we can estimate CATE from observational or experimental data by modeling the outcome response surface.
Classical approaches to heterogeneous treatment effect (HTE) estimation typically rely on subgroup analysis defined by observed covariates. Regression-based methods require pre-specifying interactions or assume homogeneity within subgroups, which can limit their ability to capture complex patterns. Recent work therefore focuses on flexible machine learning methods that discover heterogeneity in a data-driven manner. For example, Athey and Imbens (2016) introduced causal trees to identify subpopulations with different treatment effects, and Wager and Athey (2018) extended this framework to causal forests, providing individual-level treatment effect estimates along with measures of uncertainty. Propensity score–based approaches Rosenbaum and Rubin (1983) can also be adapted for HTE analysis, though subclassification may struggle in high dimensions; methods such as BART (Hill, 2011) and other flexible learners are commonly used to mitigate these issues. In our analysis, we estimate HTEs using causal forests.
3.3.1 Causal Forests
Causal forests (Athey et al., 2019; Athey and Wager, 2019; Wager and Athey, 2018) are flexible, nonparametric methods for estimating conditional average treatment effects (CATEs) and exploring treatment effect heterogeneity across covariates. They extend the random forest framework (Breiman, 2001) by adapting both the splitting rules and estimation strategies to the causal inference setting. The procedure begins with an orthogonalization step, inspired by the double machine learning framework (Chernozhukov et al., 2018) in which both the outcome and treatment variables are residualized with respect to observed covariates. This removes variation explained by potential confounders, ensuring that differences in estimated treatment effects are not driven by baseline imbalances. Residualized variables are then used to grow decision trees where the splitting criterion is chosen to maximize heterogeneity in treatment effects between child nodes, rather than minimizing prediction error of the outcome. This design allows the trees to uncover complex, nonlinear patterns in how treatment effects vary across the population.
To prevent overfitting and obtain valid inference, causal forests employ an “honest” estimation strategy: the data are split into two parts, with one subsample used to determine the structure of the tree (i.e., the splits) and the other used to estimate treatment effects within the resulting leaves. This honest splitting ensures that the estimated treatment effects are asymptotically unbiased. Many such trees are grown on bootstrap samples of the data, and their predictions are aggregated to form the final CATE estimate for each observation, reducing variance and improving stability. Because orthogonalization reduces confounding and the splitting rule targets heterogeneity, causal forests can uncover treatment effect patterns that parametric models may miss. They also handle high-dimensional covariates and complex interactions while providing valid inference. These properties make causal forests well-suited for identifying subpopulations with differing treatment responses.
4 Results
In this section, we present the results of estimating the overall causal effect of the TFO strategy using AIPW. We then examine potential heterogeneity in treatment effects by employing the causal forest framework. We perform all calculations in R (R Core Team, 2024). We implement the AIPW estimator using the CausalGAM package (Glynn and Quinn, 2017). We fit the propensity score model using a generalized additive model for a binomial outcome with a probit link function. We fit the outcome models using a generalized additive model for a conditionally Gaussian outcome with an identity link function. We use an empirical sandwich estimator of the sampling variance from Lunceford and Davidian (2004) to conduct inference.
4.1 Average Treatment Effect
We compute the ATE separately for each season, as well as for the combined dataset that includes both seasons, using the methods described in Section 3.2. Estimating the ATE separately provides insight into the stability of the treatment effect across seasons, while combining the data allows for a larger sample size, leading to narrower confidence intervals. For both the propensity score and outcome models we include all the covariates in Table 1. Additionally, for the combined data set we include season as a covariate.
4.1.1 Checks for Assumptions
As mentioned in Sections 3.1 and 3.2, AIPW relies on conditional exchangeability in order to create balanced groups. Figure 3 is a Love plot created in R for the combined sample using the cobalt package (Greifer, 2024), which compares the standardized mean differences between the treatment and control groups of the various covariates before and after AIPW. In the original sample, eight of the covariates have standardized differences that fall outside of the rule of thumb threshold of 0.05, including the time left variable with a value of 0.74. This indicates that the groups of TFO attempts and non-attempts vary considerably, particularly in the amount of time left in the period. This is an expected feature of the data and highlights why causal methods are necessary.
Figure 3: Standardized mean differences (treatment-control) of the covariates before and after AIPW. The dotted line represents the rule of thumb threshold of 0.05.
Conversely, after weighting the observations, the standardized mean differences are much closer to 0, with all covariates falling inside the threshold of 0.05. Love plots for the individual seasons (not pictured) show a similar pattern to the combined data set. Thus the weighting procedure has created groups that are comparable in terms of the measured confounders, providing evidence that the conditional exchangeability assumption is reasonable assuming the causal model represented by our DAG. In Section 4.1.3 we address the possibility of unmeasured confounders with a sensitivity analysis.
The positivity assumption can also be assessed after fitting the AIPW model. Here we do so by looking at the distribution of the estimated propensity scores. Figure 4 gives histograms of the propensity scores for both groups. Other than perhaps a couple of control observations with the smallest estimated propensity scores, there is complete overlap between the two distributions. Thus we can say that the positivity assumption is reasonable for the combined data set. Similar plots for the individual seasons (not pictured) reveal that the positivity assumption is reasonable in those analyses as well.
Figure 4: Density of the estimated propensity scores for treatment and control groups.
4.1.2 Treatment Effect Estimates
Season Average Treatment Effect
Estimate 95% Confidence Interval p-value
18-19 0.55 (0.31, 0.78) <0.001<0.001
21-22 0.77 (0.53, 1.00) <0.001<0.001
Both 0.66 (0.49, 0.83) <0.001<0.001
Table 3: ATE estimates, 95% confidence intervals, and p-values for a test of a non-zero effect based on AIPW for both the individual and combined seasons.
Table 3 gives the treatment effect estimates and 95% confidence intervals for both individual seasons and the combined data set. We estimate that if teams attempted all their TFO opportunities their average POD would be about 0.66 points higher than if they did not attempt any of their TFO opportunities, based on the combined data. A 95% confidence interval for this effect is (0.49, 0.83). The 2021-2022 ATE estimate of 0.77 is slightly higher than the 2018-2019 estimate of 0.55, but there is considerable overlap in their confidence intervals, suggesting that the variation across individual seasons may not be significant. Overall, the positive ATE across both seasons indicates that the TFO strategy consistently increases the points differential by the end of the period. The slightly more than half a point gain may seem like a minor advantage, but NBA teams are often willing to spend effort and resources in pursuit of even relatively small edges.
4.1.3 Sensitivity to Unmeasured Confounding
While we have utilized a reasonable set of covariates based on the available data, there exists the possibility of unmeasured confounding, a challenge to many causal analyses. It is natural to ask the question of how different our estimated treatment effects might be in the presence of such unmeasured confounding. Zhao et al. (2019) utilize a marginal sensitivity model to construct bootstrap confidence intervals for AIPW estimators based on a parameter $\Lambda$ that measures the degree of unmeasured confounding. Specifically, in the case of a single binary unmeasured confounder $U$, the odds of an individual with $U=1$ being in the treatment group is $\Lambda$ times higher than the odds of an individual with $U=0$ being in the treatment group. As a concrete example, suppose that immediately upon his team gaining an opportunity, a coach gives a recommendation to his players about whether or not to attempt the TFO. This could certainly have an effect on whether or not the players actually attempt the TFO. Suppose that situations in which the coach recommends the attempt have $\Lambda$ times higher odds to end up as an attempt than situations in which the coach does not recommend an attempt.
Figure 5 gives 95% confidence intervals for the ATE of a TFO attempt on the POD from the combined data set as a function of $\Lambda$ values ranging from 1.05 to 1.5, calculated using the bootsens R package (Zhao, 2018). Note that the intervals are entirely above 0 for values of $\Lambda$ up to 1.35, but contain 0 for $\Lambda\geq 1.4$ Using our example, this would mean that even if the odds of attempting a TFO increase by up to 35% when a coach recommends an attempt, the treatment effect would still be statistically significant for a two-sided test at a significance level of $\alpha=.05$. However, if those odds increase by more than 40%, the effect would no longer be statistically significant. For the 2018-2019 and 2021-2022 seasons, respectively, the treatment effect would still be statistically significant for $\Lambda$ values of up to 1.2 and 1.4, respectively.
Figure 5: 95% confidence intervals for the ATE given varying degrees of unmeasured confounding indicated by $\Lambda$.
For a more concrete example, consider the TFO attempt in the second period of the game between the Warriors and the Suns mentioned in Section 2.3. Our model estimated the propensity score for that TFO opportunity at 0.6654, which would imply the odds of attempting the TFO were around 2. If the true propensity score, given an unmeasured confounder, was 0.7, the odds would be 2.33 which would be 1.17 times higher than our estimate. However, if the true propensity score, given an unmeasured confounder, was 0.75, the odds would be 3 which would be 1.5 times higher than our estimate. In both individual seasons as well as the combined season, an overall $\Lambda=1.17$ would not be enough to change our conclusion, but an overall $\Lambda=2$ would.
Thus our results are somewhat sensitive to unmeasured confounding, particularly for the 2018-2019 season. However, in this regard having a treatment definition that relies on actions, rather than intent, may be beneficial, as it may lessen the strength of any potential confounding relationship. In our hypothetical example, due to the complexities of an NBA possession and the general autonomy of NBA players, the percentage of attempts out of all TFO opportunities may not be substantially different based on the coach’s recommendation or other unmeasured confounders.
4.1.4 Sensitivity to Treatment Definition
As mentioned in Section 2.1, another potential concern with our analysis is the somewhat arbitrary nature of our treatment definition. We chose this definition based on the structure of NBA games, before looking at any data. However, we allow that other analysts or NBA personnel might argue for slightly different cutoffs in the TFO definition. To investigate the robustness of our treatment effect estimate to these differences, we estimate the ATE of the TFO strategy for the combined data set, along with 95% confidence intervals, for a variety of other possible cutoffs. These results are presented in Figure 6. The horizontal axis represents the beginning and end of the TFO opportunity window, while the panels represent the cutoff for a TFO attempt. Thus the left-most category, 38-30, in the bottom panel represents a definition in which a team needs to gain possession with between 38 and 30 seconds left in the period to gain a TFO opportunity, and must take a shot or get fouled with at least 29 seconds left in the period for the opportunity to be considered an attempt. The vertical axis is the estimated ATE. The red circle represents the original definition with cutoffs of 43, 35, and 28 seconds, respectively.
Figure 6: Treatment effect estimates and 95% confidence intervals for alternative treatment definitions.
The treatment effect estimate using our original definition is the largest among the combinations of cutoffs we tested. However, there is considerable overlap between the original confidence interval and those with cutoffs fairly similar to our original definition. Note that definitions represented by the extreme left and right sides of the figure are, in our opinion, much less realistic and much less likely to be supported by other analysts or NBA personnel. For a broad swath of reasonable cutoffs, the estimated treatment effect is between 0.4 and 0.6. Thus while the treatment effect may be slightly higher with our definition than with others, the practical impact on policy may be minimal.
4.1.5 Assessing Fourth Period Data Exclusion
In Section 2.2, we motivated the exclusion of fourth period and overtime possessions due to end-of-game strategic differences. This subsection provides empirical evidence supporting that choice. Table 4 reports average treatment effect estimates by game period for the 2021–2022 season. The first three periods yield similar estimates, whereas the estimate for the fourth period and overtime is substantially attenuated and statistically indistinguishable from zero. This pattern suggests that end-of-game possessions follow meaningfully different strategic dynamics than those in earlier periods, justifying their exclusion from our main analysis.
Period Average Treatment Effect
Estimate 95% Confidence Interval p-value
1 0.81 (0.39, 1.23) <0.001<0.001
2 0.76 (0.31, 1.20) <0.001<0.001
3 0.60 (0.20, 1.00) $0.004$
4/OT 0.12 (-0.22, 0.46) $0.463$
Table 4: ATE estimates, 95% confidence intervals, and p-values for a test of a non-zero effect based on AIPW for the 2021-2022 season. These effects are conditional on the period, with the fourth period and overtime periods combined.
4.2 Heterogeneous Treatment Effect
The average treatment effect indicates that the TFO strategy is beneficial on average, but its value may vary across game contexts. To explore when the strategy may be more or less effective, we examine heterogeneous treatment effects (HTE) using causal forests.
We augment the dataset with three additional variables—the differences in the teams’ maximum and median ratings, and the absolute score margin—which capture relative team strength and game competitiveness. We first fit a causal forest using all variables, compute variable importance, and retain those accounting for 95% of total importance (excluding only period indicators and absolute score margin). As a check, the causal-forest ATE for the combined data set is 0.61 (SE 0.11), consistent with AIPW estimates.
4.2.1 Overall Heterogeneity
To begin our analysis, we first assess whether there is any evidence of treatment effect heterogeneity in the data. This is done through two separate tests.
First, the test calibration procedure evaluates the null of no treatment effect heterogeneity. Once the causal forest is trained, heterogeneity is assessed via a Calibration Test (Chernozhukov et al., 2018), which regresses the outcome on the mean forest prediction (ATE) and the out-of-bag differential forest prediction (CATE). As shown in Table 5, the differential forest prediction yields a p-value of 0.17, providing no significant evidence against the null.
Estimate Std Error t Value p Value
Mean Forest Prediction 1.00511 0.17876 5.6227 <0.001<0.001
Differential Forest Prediction 0.63352 0.67520 0.9383 0.1741
Table 5: Summary of Test Calibration. The test for heterogeneity in treatment effects, based on the differential forest prediction, yields a p- value greater than 0.1. This indicates insufficient evidence to reject the null hypothesis of no heterogeneity in treatment effects.
As a complementary approach, we also compute the RATE statistic (Yadlowsky et al., 2025) RATE evaluates whether a prioritization rule based on predicted CATEs improves outcomes when treatment is selectively applied. We train the causal forest on 2018 data, predict CATEs for 2021, and then form the prioritization rule. The resulting RATE estimate is 0.0408 with a standard error of 0.1176, obtained via bootstrap (Efron, 1982) with 1000 samples. The confidence interval includes zero, indicating that the prioritization rule does not identify subgroups with reliably higher gains from treatment. This aligns with the test calibration results, reinforcing the conclusion that we lack strong statistical evidence of heterogeneity.
4.2.2 Heterogeneity Across Variables
Although global tests suggest limited evidence of heterogeneity, it is still possible that treatment effects vary systematically with respect to specific variables. Prior work (De Bosscher et al., 2009) highlights settings in which overall tests fail to detect heterogeneity, yet meaningful variation appears along particular covariates. Motivated by this, we examine two measures of relative team strength: the difference in maximum player ratings between teams and the difference in median player ratings. These variables are natural candidates for effect moderation, as differences in team quality could plausibly influence the payoff of attempting a TFO.
We again rely on the RATE framework, constructing prioritization rules based solely on each variable’s value. This allows us to test whether games featuring larger rating differences exhibit systematically different treatment effects. Table 6 summarizes the RATE estimates. Both are small in magnitude, with standard errors of comparable size, and neither is statistically significant. This suggests that even along these targeted axes—where we might reasonably expect strategic differences—the data do not provide evidence that the effectiveness of the TFO strategy changes meaningfully. Thus, we conclude that there is no substantial evidence for heterogeneity in the treatment effect across the investigated variables.
Estimate Standard Error
Rating Max Diff 0.028 0.085
Rating Mean Diff 0.0825 0.087
Table 6: RATE estimates for investigating treatment effect heterogeneity across selected variables. The lack of statistical significance indicates minimal evidence of heterogeneity.
5 Discussion
This study explores the effectiveness of the two-for-one strategy in the NBA, converting a real-world sports strategy into a mathematical framework to enable causal inference. Using play-by-play data from the 2018-19 and 2021-22 NBA seasons, we analyze the impact of implementing this strategy on game outcomes by estimating the average increase in the point differential if all teams presented with a TFO opportunity attempt them versus if they do not attempt them. Our findings indicate that attempting the TFO has a statistically significant positive effect, suggesting it can be advantageous for teams in real game scenarios.
The results also consider potential heterogeneity in the effectiveness of the two- for-one strategy using the causal forest framework. While we explore overall heterogeneity and heterogeneity across two selected variables, our results do not provide significant evidence of differential effects across subsets of the data. This lack of observed heterogeneity may imply that the strategy’s impact is relatively uniform across various game contexts and player configurations, though additional factors such as small sample size may also limit our ability to detect significant variation.
While these findings contribute valuable insights, there are limitations to our analysis. First, the dataset lacks spatial information regarding players’ positions on the court during each play, which could refine estimates of shot difficulty and improve causal estimates. Such information would enhance the understanding of strategic choices made by players and coaches, but access is constrained as spatial tracking data is often proprietary. Additionally, our insignificance in heterogeneity estimates could result from the limited scope of our dataset, which includes only two seasons. Repeating the analysis over a larger dataset, potentially spanning ten seasons or more, could provide deeper insights and improve the statistical power of our heterogeneity analysis.
Another natural extension would be to consider leagues other than the NBA, such as the Women’s National Basketball Association (WNBA) or the National Collegiate Athletic Association (NCAA). Play-by-play information is generally available for these organizations as well, and it would be informative to see if similar treatment effects are found for the TFO strategy there.
In conclusion, while our study demonstrates the positive impact of the TFO strategy in NBA games, future work with extended datasets and enhanced play-by-play information could further validate and expand on these findings. By overcoming data limitations and increasing sample size, subsequent research may uncover new dimensions of heterogeneity and provide a more comprehensive understanding of how various contextual factors influence the efficacy of this strategy.